<e> Show that if P and Q are vector Subspaces of a vector space V then P ∩ Q is also a vector subspace of V.


1
Expert's answer
2022-02-20T17:28:06-0500

Define "x,y\in P\cap Q" and also "\alpha \in \mathbb{F}".

we have to prove that "P\cap Q\neq \phi" "\alpha(x+y) \in P\cap Q"

"\text{since P and Q are subspaces of V THEN , P }≠\phi \text{ and } Q≠\phi\implies P\cap Q\neq \phi"

"\text{since P and Q are subspaces of V }\implies \alpha(x+y) \in P \text{ and } \alpha(x+y) \in Q"

Hence, "\alpha(x+y) \in P\cap Q"

Thus, we may conclude that, "P\cap Q" is a subspace of "V"